# Unification of Exceptional Points and Transmission Peak Degeneracies in a Highly Tunable Magnon-Photon Dimer

### v2

@@ Abstract
Exceptional points (EPs), spectral singularities arising in non-Hermitian dynamical systems, have drawn widespread attention for their potential to enhance sensor capability by leveraging characteristic square-root frequency splitting. However, the advantages of EP sensors have remained highly contested, primarily due to their inherent hypersensitivity to model errors and loss of precision arising from eigenbasis collapse, quantified by the Petermann noise factor. Recently, it has been suggested that practical sensor implementations should instead utilize transmission peak degeneracies (TPDs), which exhibit square-root transmission splitting while maintaining a complete eigenbasis, potentially mitigating the Petermann noise around EPs. This work presents a microwave magnon-photon dimer with near-universal tunability across all relevant parameters, which we use to investigate various TPDs experimentally. We demonstrate that two-dimensional EP and TPD configurations based on coupled oscillators can be unified into a general theory and experimental framework utilizing synthetic gauge fields, which we present and validate through careful experiments on two representative TPDs. We also introduce practical metrics to quantify the performance and robustness of TPDs beyond just the Petermann noise factor. Our formalism and experimental methods unify previous EP and TPD configurations and may lead to the realization of robust TPD-enhanced sensors.
@@
I Introduction
Hermitian systems, defined by self-adjoint operators $H=H^{\dagger}$, exhibit conservative dynamics with purely real eigenvalues that ensure energy conservation and time-reversible dynamics [1]. In contrast, non-Hermitian dynamical systems, often arising from asymmetric dissipation, gain, or non-reciprocal interactions, are governed by non-self-adjoint operators, $H\neq H^{\dagger}$ with complex eigenvalues and do not conserve energy [2]. Non-Hermitian Hamiltonians have been extensively studied as models for open systems, including atoms in external fields, microwave cavities, and qubits coupled to oscillators or spin baths [3, 4, 5]. A key feature of non-Hermitian systems is their spectral degeneracies. An exceptional point (EP) occurs when two or more eigenvalues and their eigenvectors coalesce [6, 7]. Unlike Hermitian degeneracies (called diabolic points, or DPs), where only eigenvalues coalesce, EPs represent critical singularities in parameter space and exhibit non-analytic and topological spectral behavior [8, 9, 10].
Due to their fascinating characteristics, EPs have been engineered in a variety of systems, most commonly with balanced loss and gain in $\mathcal{PT}$-symmetric configurations, where EPs arise from a breaking of $\mathcal{PT}$-symmetry [11, 12, 13, 14, 15, 16, 17, 18, 19]. While probing EPs in open quantum systems typically requires cryogenic temperatures, semiclassical platforms, such as coupled magnon modes operating at room temperature, also exhibit rich behavior, such as enhanced magnonic frequency combs [20].
Over the past decade, numerous proposals and experiments have investigated the potential of EPs to enhance sensing [21, 22, 23, 24, 25]. Near an EP, eigenvalues exhibit a characteristic square-root response when subjected to perturbations in specific system parameters [16]. Despite their promise and early demonstrations [26], practical implementations of EP sensors have encountered considerable technical hurdles. Specifically, EP-based sensors require precise, isolated parameter tuning in high-dimensional parameter spaces, where even slight deviations due to noise or model errors rapidly diminish the enhanced response. This parameter hypersensitivity poses a significant, but surmountable, engineering challenge. More fundamentally, however, the eigenvector collapse at EPs—responsible for their characteristic square-root sensitivity enhancement—inherently enhances noise. This intrinsic noise enhancement at EPs, quantified by the divergence of the Petermann noise factor [27, 28], fundamentally challenges achievable improvements in signal-to-noise ratio for EP-based sensors [29]. Consequently, the viability of frequency-splitting-based EP sensors has been the subject of contentious debate [30, 31, 27, 32, 33, 34, 35].
@@ I Introduction
Previous experimental platforms for studying EPs and TPDs most commonly include microcavities operating in the optical regime [38, 25, 26, 39], electrical circuits in the RF regime [25, 37, 40], or cavity magnonics in the microwave regime [41, 42]. Cavity magnonics exploits strong magnon-photon coupling to enable quantum sensing by hybridizing spin excitations with electromagnetic fields [43]. Crucially, the coupling phase between magnonic and photonic modes acts as an additional control knob that can be used to alter the coupling structure and modify the system dynamics [44, 45]. Specifically, phases play a crucial role in “loop-coupled” geometries, where they combine to form effective U(1) gauge fields [46] and can be used to realize tunable coherent and dissipative couplings [47]. These synthetic gauge fields in cavity magnonics can also enable non-reciprocal interactions by breaking symmetries, as demonstrated in a generalized Hatano-Nelson model [48, 49].
This work introduces a flexible theoretical and experimental framework that unifies two-dimensional coupled-oscillator-based EP and TPD configurations, establishing a foundation for future TPD-based sensors to fulfill the original promise of EP-enhanced sensing. We present a tunable microwave magnon-photon dimer with independent control of resonance frequencies, dissipation rates, coupling strength, and coupling phase. We use these control knobs to survey a two-dimensional surface of EPs and TPDs embedded within the six-dimensional parameter space, traversed using a synthetic gauge field on the coupling term, $Je^{i\phi}$, effectively forming a non-Hermitian beam splitter [50]. We demonstrate two sets of EPs and TPDs and study how the TPDs move with parameter choice to locations with a complete—or even orthogonal—eigenbasis. Our platform flexibility may enable the optimization of sensor characteristics based on robustness and performance metrics introduced in Sec. IV.0.3. The formalism, platform, and experimental methods presented here are versatile testbeds for TPD sensors and general non-Hermitian physical phenomena.
II Experimental Architecture
@@
Our experimental architecture is a coupled magnon-photon system, forming a dimer of dissipative harmonic oscillators, each characterized by a resonant frequency $f$ and total dissipation rate $\kappa$. The photon mode is the lowest harmonic of a three-dimensional microwave cavity [51], while the magnon mode is realized using a 1 mm YIG sphere. YIG exhibits a narrow ferrimagnetic resonance linewidth and is a widely used material in cavity magnonics [43], making it an ideal candidate for our realization of an independent magnon mode. When biased with a static magnetic field, the electron spins in YIG collectively precess about the field axis, forming a resonance analogous to a ferromagnetic resonance (FMR) (for more details, see Supplementary Information). The Larmor frequency depends linearly on the applied field as $f_{y}=\mu_{B}B_{z}/2\pi$, where $\mu_{B}$ is the Bohr magneton, consistent with the simplified Kittel formula for a uniformly magnetized sphere [52, 53]. The magnon mode is coupled to using two orthogonal wire bonds that encircle the YIG sphere. Without the YIG, the loops are electromagnetically isolated, preventing power transmission between them. When the YIG is inserted, its spin resonance enables microwave coupling between the input and output loops only at frequencies near the Larmor frequency. Therefore, the oscillator hosting the YIG passes signals around $f_{y}$, with a FWHM $\kappa_{y}$ while attenuating others [54].
@@
The magnon and photon modes are spatially separated by approximately one meter using SMA microwave coaxial lines, as shown in Fig. 1a. By utilizing unidirectional microwave amplifiers, the coupling between the modes can be split into independent coupling paths, each of which consists of a fixed-gain amplifier followed by a digital attenuator, effectively forming a digitally controlled variable-gain amplifier with gain $G$ that we use to tune the overall coupling $J$ [55]. To ensure symmetric coupling rates, we carefully measure the insertion losses of components on each coupling path and offset the digital attenuators accordingly. The coupling phase $\phi$ introduces phase non-reciprocity in the coupling, which, while not necessarily breaking reciprocity in the strict electromagnetic sense, produces direction-dependent phase evolution, effectively producing a synthetic gauge field [46]. To tune the coupling phase, we insert a digital phase shifter in one of the coupling paths (Fig. 1a). The phase shifter enables continuous tuning of the complex coupling coefficient, $Je^{i\phi}$, across the full range $\phi\in[0,2\pi)$ while preserving the coupling magnitude. We focus on two representative cases: reciprocal phase (RP) coupling ($\phi=0$) and anti-reciprocal phase (ARP) coupling ($\phi=\pi$). These two phase configurations produce distinct coupling dynamics between the magnon and photon modes, consistent with traditional level repulsion and level attraction [44]. The system exhibits level repulsion for RP coupling, characterized by an avoided crossing whose gap increases with the coupling strength $J$ (Fig. 1b,c,f,g). Modifying the average dissipation rate,
$\bar{\kappa}=(\kappa_{c}+\kappa_{y})/2,$ (1)
changes the linewidths of the hybridized modes while preserving the repulsion behavior (Fig. 1f,g). In contrast, ARP coupling results in level attraction, where the hybridized modes coalesce and split as the individual frequencies are detuned. The extent of the coalescence region increases with $J$ (Fig. 1d,e).
To achieve full parameter tunability, each mode is equipped with a self-feedback loop consisting of a phase shifter with phase $\psi_{c,y}$ and an effective variable-gain amplifier with gain $g_{c,y}$ (Fig. 1a), where the subscripts $(c,y)$ correspond to the cavity and YIG self-feedback loops, respectively. Specifically, $g_{c}$ is controlled using a coarse digital attenuator, while $g_{y}$ is controlled using both a digital attenuator and an analog attenuator in series, for higher resolution control. $\psi_{c,y}$ and $g_{c,y}$ enable tunable constructive and destructive feedback interference, adding additional control over $f$ and $\kappa$ for each mode. The interactions between $\psi_{c,y},g_{c,y}$ and the mode parameters are not explicitly modeled; instead, we utilize $\psi_{c,y},g_{c,y}$ to achieve the desired $(f_{c},f_{y})$ and $(\kappa_{c},\kappa_{y})$. At each step during the experiment, we measure and extract fitted values of $f_{c,y},\kappa_{c,y}$, and their uncertainties $\sigma(f_{c,y}),\sigma(\kappa_{c,y})$. To achieve this, the dimer includes microwave switches to enable three distinct operating modes: cavity readout, YIG readout, and normal operation. In cavity (YIG) readout operation mode, the magnon-photon coupling is disabled, and transmission is measured through the cavity (YIG) mode in isolation, allowing for measurement of $f_{c,y},\kappa_{c,y}$ and their associated uncertainties $\sigma({f_{c,y}}),\sigma(\kappa_{c,y})$. During normal operation, magnon-photon coupling is enabled, the cavity is driven, and the YIG is measured. Figure 1 illustrates the complete experimental schematic (Fig. 1a and associated transmission measurements in normal operation (Fig. 1b–g), highlighting the independent tunability of all model parameters.
III Theory
One of the core contributions of this work is a unified state-space framework for analyzing EPs and TPDs in dimers. Our state-space model is algebraically equivalent to the standard coupled-mode or quantum-Langevin formalisms at the linear, semiclassical level, yet offers practical advantages for our study. Specifically, a single $(\bm{A},\bm{B},\bm{C})$ representation captures both the eigenspectrum that defines EPs and the transmission spectrum that reveals TPDs. The state space model matches all the relevant features in the experimental data: peak locations, TPDs, EPs, and instability transitions. Our theoretical model is motivated by the dissipative Hamiltonian of magnon-photon coupling studied in cavity magnonics, where a coupling phase in one of the off-diagonal elements induces phase non-reciprocity [47]. For completeness, in the Supplementary Information, we motivate this model using an effective quantum Hamiltonian with non-reciprocal hopping between modes and the corresponding semi-classical limit, and recover the state-space matrices $(\bm{A},\bm{B},\bm{C})$.
@@ III Theory
The dynamics of the dimer are governed by semiclassical equations of motion
$\dot{\bm{\alpha}}$ $=$ $\bm{A}\bm{\alpha}+\bm{B}u,$ (2)
where $\bm{\alpha}=[\alpha,m]^{\mathrm{T}}$ denotes the complex coherent state amplitudes of the photon ($\alpha$) and magnon ($m$) modes. $\bm{B}\in\{[1,0]^{\mathrm{T}},[0,1]^{\mathrm{T}}\}$ is the drive vector, where $\bm{B}=[1,0]^{\mathrm{T}}$ ($\bm{B}=[0,1]^{\mathrm{T}}$) represents driving the cavity (YIG) with drive strength $u$. Complex mode amplitudes are measured by $\bm{C}\bm{\alpha}$ with $\bm{C}\in\{[1,0],[0,1]\}$ the readout vector, $\bm{C}=[1,0]$ ($\bm{C}=[0,1]$) corresponds to reading out the cavity (YIG). The theoretical model in Eq. 2 reproduces relative power between the two modes rather than the absolute microwave power in dBm, which is sufficient for analyzing TPDs, EPs, frequency-peak-splitting, and instability.
In our architecture, $f_{c,y},\kappa_{c,y}$ are all independently tunable. However, $f_{y}$ and $\kappa_{y}$ are controlled with high-resolution analog voltage sources, while $f_{c}$ and $\kappa_{c}$ rely on discrete digital control with coarser resolution. Thus, to enable fine-grained experimental sweeps, we fix $f_{c}$ and $\kappa_{c}$, and sweep $f_{y}$ and $\kappa_{y}$. Accordingly, it will be convenient to consider parameter detunings defined by
$\Delta_{f}$ $\equiv f_{c}-f_{y},$ (3)
$\Delta_{\kappa}$ $\equiv\frac{\kappa_{c}}{2}-\frac{\kappa_{y}}{2}.$ (4)
Utilizing the relative detuning parameters defined in Eqs. 3 and 4, we construct a semi-classical dynamical matrix in the rotating frame of the drive at frequency $f_{d}$ (see Supplementary Information for derivation)
@@
$\bm{A}=\begin{bmatrix}-\mathrm{i}(f_{c}-f_{d})-\frac{\kappa_{c}}{2}&-\mathrm{i}J\\ -\mathrm{i}Je^{i\phi}&-\mathrm{i}(f_{c}-f_{d}-\Delta_{f})+(\Delta_{\kappa}-\frac{\kappa_{c}}{2})\end{bmatrix}$ (5)
Although the expression in Eq. 5 includes both relative parameters $\Delta_{\kappa},\Delta_{f}$ and absolute parameters $f_{c}$, $\kappa_{c}$, this representation reduces the complexity of the expressions that follow and reflects a more natural coordinate system for analyzing the dynamical behavior of the system. Furthermore, although we use dimensional parameters $\Delta_{f},\Delta_{\kappa}$ here, the theory can also be nondimensionalized by normalizing all parameters by $J$ (see Supplementary Information).
Next, we consider the eigenspectrum of $\bm{A}$, which we use to parameterize EPs in the system. Following that, in Sec. III.0.2, we consider the transmission spectrum, used to parameterize TPDs. Finally, in Sec. III.0.3, we summarize and unify the results of EPs, TPDs, and instability transition locations in the system. The discrepancy between the EP and TPD for the RP configuration [36, 37] and ARP configuration [25] have been explored separately and experimentally validated in previous work. Here, we generalize this analysis to all values of the coupling phase $\phi$, and present the representative cases $\phi=0$ and $\phi=\pi$. We extend this to $\phi=\pi/2$ in the Supplementary Information.
III.0.1 Eigenspectrum Analysis
The eigenvalues of $\bm{A}$ in Eq. 5, denoted $\lambda_{\pm}$, encode information about the physical characteristics of the system dynamics. Here, the real part of each eigenvalue corresponds to dissipation ($\text{Re}(\lambda)<0$) or gain ($\text{Re}(\lambda)>0$), and the imaginary part corresponds to the oscillation frequency of the modes. This convention differs from the more typical formulation in coupled-mode theory, where the system Hamiltonian $H$ generates dynamics through $\mathrm{i}H$, leading to the opposite association between real and imaginary parts of the spectrum [56]. Our convention reflects the semiclassical treatment of driven-dissipative systems, consistent with standard practice in input–output theory for non-Hermitian photonics analyzed in state-space form [57, 55].
The eigenvalues $\lambda_{\pm}$ can be written as a symmetric pair around a central complex value $\lambda_{0}$, with splitting given by $\Delta_{\lambda}=\lambda_{+}-\lambda_{-}$, or in terms of the model parameters as

### v3

@@ Abstract
Exceptional points (EPs), spectral singularities arising in non-Hermitian dynamical systems, have drawn widespread attention for their potential to enhance sensor capability by leveraging characteristic square-root frequency splitting. However, the advantages of EP sensors have remained highly contested, primarily due to their inherent hypersensitivity to fabrication errors and loss of precision arising from eigenbasis collapse, quantified by the Petermann noise factor. Recently, it has been suggested that practical sensor implementations should instead utilize transmission peak degeneracies (TPDs), which exhibit square-root transmission splitting while maintaining a complete eigenbasis, avoiding the Petermann noise around EPs. This work presents a microwave magnon-photon dimer with near-universal tunability across all relevant parameters, which we use to survey a surface of EPs and TPDs experimentally, including a newly introduced phase-non-reciprocal TPD. We demonstrate that two-dimensional EP and TPD configurations based on coupled oscillators can be unified into a general theory and experimental framework utilizing synthetic gauge fields, which we present and validate through careful experiments on six representative TPDs. We also introduce practical figures of merit to quantify the performance and robustness of TPDs beyond just the Petermann noise factor. Our formalism and experimental methods unify previous EP and TPD configurations, and may lead to the widespread realization of robust TPD-enhanced sensors.
@@
Over the past decade, numerous proposals and experiments have investigated the potential of EPs to enhance sensing [1, 2, 3, 4, 5, 6]. Near an EP, eigenvalues exhibit a characteristic square-root response when subjected to perturbations in specific system parameters [7]. Despite their promise and early demonstrations [8], practical implementations of EP sensors have encountered considerable technical hurdles. Specifically, EP-based sensors require precise, isolated parameter tuning in high-dimensional parameter spaces, where even slight deviations due to fabrication errors rapidly diminish the enhanced response, a phenomena known as parameter hypersensitivity. More fundamentally, however, the eigenvector collapse at EPs—responsible for their characteristic square-root response—inherently enhances noise. This intrinsic noise enhancement at EPs, quantified by the divergence of the Petermann noise factor [9, 10], fundamentally challenges achievable improvements in signal-to-noise ratio for EP-based sensors [11]. Consequently, the viability of frequency-splitting-based EP sensors has been the subject of contentious debate [12, 13, 9, 14, 15, 16, 17].
@@
Previous experimental platforms for studying EPs and TPDs most commonly include microcavities operating in the optical regime [20, 8, 21, 22], electrical circuits, including superconducting circuits, in the RF regime [5, 19, 23, 24, 25, 26], or cavity magnonics in the microwave regime [27, 28]. Cavity magnonics exploits strong magnon-photon coupling to enable quantum sensing by hybridizing spin excitations with electromagnetic fields [29]. Crucially, the coupling phase between magnonic and photonic modes acts as an additional control knob that can be used to alter the coupling structure and modify the system dynamics [30, 31, 32]. Specifically, phases play a crucial role in “loop-coupled” geometries, where they combine to form effective U(1) gauge fields [33] and can be used to realize tunable coherent and dissipative couplings [34, 35]. These synthetic gauge fields in cavity magnonics can also enable non-reciprocal interactions by breaking symmetries, as demonstrated in a generalized Hatano-Nelson model [36, 37].
This work introduces a flexible theoretical and experimental framework that unifies two-dimensional coupled-oscillator-based EP and TPD configurations, establishing a foundation for future TPD-based sensors. We present a tunable microwave magnon-photon dimer with independent control of resonance frequencies, dissipation rates, coupling strengths, and coupling phase. We use these control knobs to survey a surface of EPs and TPDs embedded within the six-dimensional parameter space, traversed using a synthetic gauge field on the coupling term, $Je^{i\phi}$, effectively forming a non-Hermitian beam splitter [38]. We demonstrate three sets of EPs and TPDs and study how the TPDs move with parameter choice to locations with a complete—or even orthogonal—eigenbasis. Our platform flexibility may enable the optimization of sensor characteristics based on figures of merit introduced in Sec. III.1. The formalism, platform, and experimental methods presented here are versatile testbeds for TPD sensors and general non-Hermitian physical phenomena.
I Experimental Architecture
@@
Our experimental architecture is a coupled magnon-photon system, forming a dimer between the lowest harmonic of a three-dimensional microwave cavity [39] and a magnon mode realized using a yttrium iron garnet (YIG) sphere of radius 1 mm. In the presence of a static magnetic field, electron spins in YIG collectively precess about the field axis, forming a resonance analogous to ferromagnetic resonance (FMR) (for more details, see Supplementary Information). The Larmor frequency is $f_{y}=\mu_{B}B_{z}/2\pi$, where $\mu_{B}$ is the Bohr magneton, consistent with the simplified Kittel formula for a uniformly magnetized sphere [40, 41]. The YIG mode is coupled to using two orthogonal wire bonds that encircle the YIG sphere. Without the YIG, the loops are electromagnetically isolated, preventing power transmission between them. When the YIG is inserted, its spin resonance enables microwave coupling between the input and output loops only at frequencies near the Larmor frequency. Therefore, the oscillator hosting the YIG passes signals around $f_{y}$, with full-width-half-maximum $\kappa_{y}$, while attenuating others, manifested as a characteristic Lorentzian transmission spectra [42].
@@
The magnon and photon modes are spatially separated by approximately one meter [43]. Coupling between the modes is split into independent coupling paths using SMA microwave coaxial cables (Fig. 1a). Following previous work, we place a tunable directional amplifier in each coupling path to symmetrically tune the coupling strength $J$, and include a phase shifter $\phi$ in one arm to introduce phase non-reciprocity in the coupling [35] (Methods). This phase non-reciprocity does not necessarily break reciprocity in the strict electromagnetic sense, but produces direction-dependent phase evolution, effectively producing a synthetic gauge field [33]. The resulting complex coupling coefficient, $Je^{i\phi}$, is tunable across the full range $\phi\in[0,2\pi)$ while preserving the coupling magnitude, allowing access to distinct coupling regimes. At $\phi=0$, or reciprocal phase coupling, the system exhibits level repulsion, characterized by an avoided crossing whose gap increases with the coupling strength $J$ (Fig. 1b,c,f,g). Modifying the average dissipation rate, $\bar{\kappa}=(\kappa_{c}+\kappa_{y})/2$, modifies the linewidths of the hybridized modes while preserving the repulsion behavior (Fig. 1f,g). At $\phi=\pi$, or anti-reciprocal phase coupling, level attraction is observed, where the hybridized modes coalesce and split as the individual frequencies are detuned [30] (Fig. 1d,e). The extent of the coalescence region increases with $J$, as expected. For all other phase values, such as $\phi=\pi/2$, the two couplings still share the same magnitude but their complex phases differ by an angle that cannot be removed by a single phase redefinition or gauge transformation, resulting in true phase-non-reciprocal dynamics.
To complete the six-parameter control of the dimer, each mode is further equipped with a self-feedback loop consisting of a phase shifter with phase $\psi_{c,y}$ and an effective variable-gain amplifier with gain $g_{c,y}$ (Fig. 1a), where the subscripts $(c,y)$ correspond to the cavity and YIG self-feedback loops, respectively. Adjusting $\psi_{c,y}$ and $g_{c,y}$ provides in-situ tuning of the mode frequency $f_{c,y}$ and dissipation rate $\kappa_{c,y}$. At each step during an experiment, we probe both modes individually with coupling disabled to extract $f_{c,y}$, $\kappa_{c,y}$ and uncertainties $\sigma(f_{c,y}),\sigma(\kappa_{c,y})$ (Methods). Furthermore, any desired $f_{y},\kappa_{y}$ can be attained using our regression-based parameter control system that optimally sets $\psi_{y},g_{y}$ (Methods). Figure 1 illustrates the complete experimental schematic (Fig. 1a) and various transmission experiments highlighting parameter control (Fig. 1b–g).
II Non-Hermitian Dynamics
@@
The dynamics of the dimer are modeled using a state space model, governed by the semiclassical equations of motion
$\dot{\bm{\alpha}}$ $=\bm{\tilde{A}}\bm{\alpha}+\bm{B}u,$ (1)
$\tilde{\beta}_{\mathrm{ss}}$ $=\left|\bm{C}\bm{\tilde{A}}^{-1}\bm{B}u\right|^{2}.$ (2)
where $\bm{\tilde{A}}$ is the dynamical system matrix defined in Eq. 3, and $\bm{\alpha}=[\alpha,m]^{\mathrm{T}}$ denotes the complex coherent state amplitudes of the photon ($\alpha$) and magnon ($m$) modes. $\bm{B}\in\{[1,0]^{\mathrm{T}},[0,1]^{\mathrm{T}}\}$ is the drive vector, where $\bm{B}=[1,0]^{\mathrm{T}}$ ($\bm{B}=[0,1]^{\mathrm{T}}$) represents driving the cavity (YIG) with drive strength $u$. Complex mode amplitudes are measured by $\bm{C}\bm{\alpha}$ with $\bm{C}\in\{[1,0],[0,1]\}$ the readout vector, $\bm{C}=[1,0]$ ($\bm{C}=[0,1]$) corresponds to reading out the cavity (YIG). $\tilde{\beta}_{ss}$ in Eq. 2 represents the steady state transmission spectrum, which is only valid when all real parts of the eigenvalues of $\bm{\tilde{A}}$ are negative. When $\bm{\tilde{A}}$ becomes unstable, amplifier saturation eventually curbs the runaway gain, leading to nonlinear phenomena reported in earlier work but not considered here [35].
In the rotating frame with respect to the drive frequency $f_{d}$, the system matrix generates dynamics and is given in nondimensionalized form by
@@
$\bm{\tilde{A}}=\begin{bmatrix}-\mathrm{i}(\tilde{f}_{c}-\tilde{f}_{d})-\frac{\tilde{\kappa}_{c}}{2}&-\mathrm{i}\\ -\mathrm{i}e^{i\phi}&-\mathrm{i}(\tilde{f}_{c}-\tilde{f}_{d}-\tilde{\Delta}_{f})+\left(\tilde{\Delta}_{\kappa}-\frac{\tilde{\kappa}_{c}}{2}\right)\end{bmatrix}$ (3)
where $\tilde{f_{c}}=f_{c}/J,\tilde{\kappa}_{c}=\kappa_{c}/J$, etc., and $\tilde{\Delta}_{f}=\tilde{f}_{c}-\tilde{f}_{y}$, and $\tilde{\Delta}_{\kappa}=(\tilde{\kappa}_{c}-\tilde{\kappa}_{y})/2$ (see Supplementary Information for derivation) [32].
Figure 2: Geometry of EPs and TPDs. Parameter landscapes for $\phi=0$ (a), $\phi=\pi$ (b), and $\phi=\pi/2$ (c). Eigenspectrum characteristics include absolute imaginary eigenvalue splitting $|\text{Im}(\tilde{\Delta}_{\lambda})|$ (i), real eigenvalue splitting $\text{Re}(\tilde{\Delta}_{\lambda})$ (ii), and mean diagonal Petermann factor $\bar{K}_{2}$ (iii-iv). Crosses are EPs, where $\text{Im}(\tilde{\Delta}_{\lambda})=\text{Re}(\tilde{\Delta}_{\lambda})=0$ (i-ii), and $\bar{K}_{2}$ diverges (iii-iv) ($\bar{K}_{2}$ colorscale is clipped to 90th percentile). Transmission spectrum characteristics are determined by the contours in iii-iv, whose locations depend on $\tilde{\kappa}_{c}$. Chosen values of $\tilde{\kappa}_{c}$ match those used in experiment (Fig. 3). Cyan contours are $\mathrm{Disc}=0$ (discriminant of Eq. 9), separating the parameter space into single peak (between cyan contours horizontally) and two peak (outside cyan contours horizontally) regions. Magenta contours are $\tilde{q}=0$ (Eq. 11). We select the $\tilde{q}=0$ contour that crosses the primary EP and TPD as the independent variable in Fig. 3. Lime contours are $\text{Re}(\tilde{\Delta}_{\lambda})=(\tilde{\kappa}_{c}-\tilde{\Delta}_{\kappa})$, and separate the parameter space into stable (inferno colorscale) and unstable (gray colorscale) regimes (Eq. 2 is inapplicable in the unstable regime and not considered in this work). Circles represent TPDs associated with an EP (Primary EPs and TPDs, in red, are observed in Fig. 3), while diamonds represent TPDs not associated with an EP, referred to as rogue TPDs.
II.1 Eigenspectrum Analysis
The eigenvalues $\tilde{\lambda}_{\pm}$ of the dynamical matrix $\bm{\tilde{A}}$ are given by